# MiM: Mask in Mask Self-Supervised Pre-Training for 3D Medical Image Analysis

### v1

@@ 4.7.3 Reconstruction results
Fig. 9: Reconstruction results of the proposed method MiM on the BTCV validation dataset [43].
5 Discussion
5.1 Limitations
While our MiM framework has shown promising results, our MiM does have limitations. In our experiments, we only scaled up the MiM to a 10k pre-training dataset, which may not be sufficient to fully explore its potential. Previous works in Natural Language Processing [41] and Computer Vision [41, 68] have demonstrated that scaling up the pre-training dataset to billions of images can lead to the foundation model for their fields [41], but collecting enough high-quality 3D medical images remains challenging. Therefore, further exploration is needed to investigate the performance of MiM with a vast scale of pre-training datasets, such as 100k.
Moreover, while the hierarchical design is effective, we adopted a simple design in the experiments to consider computation costs. For the generation of multi-level volumes, we employed a random sampling strategy to select the masked patches as next-level volumes, which achieved good performance. Since not all patches are equally important [45], Wang et al. [66] proposed a hard negative mining strategy to sample the patches for reconstruction. Future exploration can design priory sampling strategy i.e., selecting the most informative patches for finer level reconstruction, to further improve the performance of MiM. Finally, as our MiM framework mainly mines the hierarchical structure for intra-volume relationships, it can be extended to explore its capacity in mining inter-volume relationships, which may further improve its performance.
5.2 Future directions and challenges
The foundation model has the potential to advance healthcare [33], and SSL is a promising direction for building the foundation model in 3D medical images [33, 1]. However, there are current limitations that need to be addressed. To build upon the 3D medical image foundation model with SSL, several promising emerging directions can be explored. (1) Building up a large-scale pre-training datasets, e.g., more than 100k volumes. (2) Exploring more downstream tasks, e.g., 3D medical image registration [86], and tumor synthesis [36]. (3) Exploring cooperation with other modalities for multi-modal pre-training, e.g., language [73, 1, 44]. Actually, (1) is the cornerstone for the foundation model in 3D medical images, and (2) can further comprehensively evaluate the performance of SSL for 3D medical image analysis. Since existing SSL methods often pre-train with a limited scale of data and evaluate on limited downstream tasks, which cannot lead to a healthcare foundation model for 3D medical images [73]. Inspired by the success of  Achiam et al. [1], Radford et al. [53], multi-modal pre-training for 3D medical image analysis also deserves more attention, which can learn robust representation by complementing the information from different modalities.
@@ 6 Conclusion
In this paper, we introduce the Mask in Mask (MiM) pre-training framework, which significantly advances 3D medical image analysis. MiM improves MAE by facilitating the learning of hierarchical visual tokens across various scales. This results in a robust representation for transfer learning in downstream application-specific tasks. By incorporating hierarchical designs, i.e., multi-level reconstruction and cross-level alignment, MiM efficiently encodes multi-granularity visual cues of structure and details into the representation. To facilitate the fair and comprehensive comparison of existing methods, we collected ten public datasets and curated two scales of pre-training datasets, i.e., 1k and 10k. Consistent with existing works, we pre-trained our MiM on a 1k dataset, and extensive experiments on thirteen datasets demonstrate MiM’s effectiveness in achieving superior performance over other methods, including segmentation and classification tasks. We also transferred the learned representation of MiM to MRI-based tasks to evaluate the cross-modality generalization ability. Additionally, we conducted an ablation study to validate the importance of each component of MiM. We conducted a hyper-parameter analysis to investigate the impact of different hyper-parameters on the performance of MiM. The results reveal that the hierarchical design of the MiM framework is crucial for achieving superior performance for 3D medical images. We further explore to scale up the pre-training dataset to 10k. The results show that the performance of MiM can be further improved by scaling up the pre-training dataset. This finding emphasizes the importance of large-scale pre-training for building the foundation model in 3D medical images. The study hopes to serve as a strong baseline for future research in 3D medical image analysis and draw more attention from the community to the importance of hierarchical design and large-scale pre-training towards the healthcare foundation model.
7 Acknowledgments
This work was supported by the Hong Kong Innovation and Technology Fund (Project No. MHP/002/22), Shenzhen Science and Technology Innovation Committee Fund (Project No. SGDX20210823103201011), and the Project of Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone (HZQB- KCZYB-2020083).

### v2

@@ 4.4.3 Reconstruction results
Figure 10: Reconstruction results on the BTCV validation dataset [53].
4.4.4 Incorporating Parameter-Efficient Fine-Tuning Methods
Parameter-efficient fine-tuning methods, such as LoRA [74] and Ladder fine-tuning [75], offer valuable insights in realistic scenarios with limited computation resources [74, 75]. Specifically, we evaluated LoRA [74] for both classification and segmentation tasks on 3D medical images using our proposed method, implementing it via the official codebase  [^1] https://github.com/microsoft/LoRA. Our implementation introduces low-rank matrices into attention and convolutional layers to approximate weight updates while freezing the original parameters during fine-tuning. As demonstrated in Table 13 and Table 14, LoRA achieves comparable performance to full fine-tuning while significantly reducing the number of updated parameters, consistent with previous findings [74] and [75]. We observed that increasing the rank $r$ improves performance, approaching that of full fine-tuning, though at the cost of increased training time and memory usage. These results show that our method can effectively leverage LoRA [74] during fine-tuning, achieving an optimal balance between computational resources and performance.
Table 13: Comparison of different LoRA fine-tuning strategies applied to our method on the MM-WHS dataset [60]. The Update Param. column shows the ratio between LoRA’s trainable parameters and the total trainable parameters, excluding the segmentation decoder which was added later for downstream tasks.
Fine-tuning strategies Update Param. (%) Memory Usage (GB) Segmentation
Train Inference DSC(%) NSD(%)
Fully fine-tuning(default) 100 16.74 0.81 91.12 96.19
LoRA [74](r=16) 10.51 17.03 0.82 91.09 96.72
LoRA [74](r=8) 5.55 16.72 0.81 90.97 96.58
LoRA [74](r=4) 2.85 16.58 0.81 91.01 96.78
Table 14: Comparison of different LoRA fine-tuning strategies applied to our method on the CC-CCII dataset [64]. The Update Param. column shows the ratio between LoRA’s trainable parameters and the total trainable parameters, excluding the classifier which was added later for classification downstream tasks.
Fine-tuning strategies Update Param. (%) Memory Usage (GB) Classification
Train Inference ACC(%) AUC
Fully fine-tuning(default) 100 14.52 0.20 94.26 99.69
LoRA [74](r=16) 11.17 17.18 0.21 94.12 98.99
LoRA [74](r=8) 8.10 16.46 0.21 93.78 98.74
LoRA [74](r=4) 5.37 16.05 0.21 93.95 98.78
@@ 5 Conclusion
In this paper, we introduce the Mask in Mask (MiM) pre-training framework, which significantly advances 3D medical image analysis. By incorporating hierarchical designs, i.e., multi-level reconstruction and cross-level alignment, MiM efficiently encodes multi-granularity visual cues of structure and details into the representation. To facilitate the fair and comprehensive comparison of existing methods, we collected ten public datasets and curated two scales of pre-training datasets, i.e., 1k and 10k. The results reveal that the hierarchical design of the MiM framework is crucial for achieving superior performance for 3D medical images. We further explore to scale up the pre-training dataset to 10k. The results show that the performance of MiM can be further improved by scaling up the pre-training dataset. This finding emphasizes the importance of large-scale pre-training for building the foundation model in 3D medical images.
Several promising directions can be explored upon this work. (1) Building up a large-scale pre-training datasets, e.g., more than 100k volumes. (2) Exploring more downstream tasks, e.g., 3D medical image registration. (4) Exploring the potential of learnable level embeddings instead of hard coding the number of level. (3) Exploring cooperation with other modalities for multi-modal pre-training, e.g., language. Actually, (1) is the cornerstone for the foundation model in 3D medical images, and (2) can further comprehend the evaluation. (3) can complement the 3D medical foundation model with the information from different modalities. (4) can further improve the method’s design.